Mortgage calculations (show all your work). If you want to buy a home for $300,000, make...

  1. Mortgage calculations (show all your work).

  1. If you want to buy a home for $300,000, make a down payment of $50,000 and take out a mortgage on the rest of the money at 6.25% compounded monthly, what will be your monthly payment to retire the mortgage in 15 years? (7 points)
  1. Consider the 60th payment. How much will the interest and principal payments be? (7 points)
Answer #1

(a) Down payment=$50000

Cost of house=$300000

Therefore loan amount,P=300000-50000=$250000

Yearly interest rate=6.25%=.0625

Monthly interest rate i=.0625÷12=.0052

No of years=15

Therefore total no of months,N=15×12=180 months

Monthly payment is given by equation,

c=rP((1+r)^N))/(((1+r)^N)-1)

=

=

=3306.661/1.543=$2143.007

(b)Principal amount at beginning of 60th payment=amount owed at end of month 59

=

=

=339497.299-147533.337=$191963.962

Therefore interest rate to be paid at 60th month=r×principal amount remaining at end of 59th month

=.0052×191963.962=$998.21

Thank you for the question.

Keep Chegging.